### v1

# A Novel Intensity-Based Criterion for Determining Exceptional Point in PT-Symmetry Coupled Array of Optical Waveguides

Abstract
In this study, we investigated the propagation pattern and the site-to-site correlation function in a PT-symmetric waveguide array with different input quantum states. Recognizing the stark difference in propagation pattern before and after the PT symmetry-breaking point, we have developed a novel, straightforward intensity-based criterion to determine the exceptional point (EP). This new criterion shows excellent agreement with those obtained by directly computing the Hamiltonian’s eigenvalues. Given the computational complexity of Hamiltonian diagonalization, our proposed criterion provides a highly efficient and valuable alternative for identifying the PT symmetry-breaking point.
Keywords: Waveguide Array, Quantum Input State, PT-Symmetry, Exceptional Point, Intensity-Based Criterion, Propagation Pattern, Site-to-Site Correlation Function.
1 Introduction
Over the past decade, discrete optical systems-particularly waveguide arrays-have played a crucial role in controlling the behavior of light [1]. Optical waveguide arrays consist of several single-mode waveguides arranged at specific distances. In this system, light is transferred between adjacent waveguides through evanescent coupling. The most effective method for manufacturing these waveguides is direct writing with femtosecond lasers [2].
The ability to control various parameters within waveguides offers a significant advantage for investigating PT-symmetric systems. PT-symmetry in quantum mechanics was first explored in 1998, following the discovery of Hamiltonians that remain invariant under both parity and time-reversal operations [3]. Notably, these non-Hermitian Hamiltonians can exhibit entirely real spectra [4]. PT-symmetry was introduced into optics due to the similarity between the paraxial wave equation and the Schrödinger equation, with the refractive index serving as the optical equivalent of the potential [5]. A PT-symmetric system in optics can be realized by adjusting the real and imaginary components of the refractive index [6]. For PT-symmetry to hold in optics, the real part of the refractive index needs to be spatially even, while its imaginary part must be odd function of position [7]. Since the imaginary part of the refractive index defines gain and loss, a PT-symmetric system can be achieved in optics by tuning these parameters accordingly. In a two-component optical system, real eigenvalues arise when the gain in one component compensates for the loss in the other. In contrast, in the weak coupling regime, if the gain from one component cannot fully compensate for the loss in the other, the system exhibits complex eigenvalues. This phase transition from real to complex spectra occurs at the EP [8].
A non-Hermitian Hamiltonian typically leads to complex eigenvalues, where the negative imaginary part represents a damping constant associated with the finite spectral linewidth of the resonance energy [9]. The real part of the eigenvalue determines this linewidth. In classical optical systems, the resonance frequencies and modes correspond to the eigenvalues and eigenstates of the Hamiltonian. Non-Hermiticity can arise due to absorption, gain, and radiation. When the Hamiltonian is non-Hermitian, i.e $\hat{H}\neq\hat{H}^{\dagger}$, its eigenstates lose their orthogonality. The non-orthogonality is maximized at the EP, where the eigenstates of the Hamiltonian coalesce, and they can no longer form a complete basis. In this scenario, the states become parallel to one another. These points are considered singularities within non-Hermitian systems. Near an EP, one or more eigenvalues and their corresponding eigenstates coalesce, leading to a phase transition. This transition, known as a PT-symmetry phase transition, results in the emergence of complex eigenvalues [8, 10].
To compute the EP, the eigenvalues of the system must be calculated, then, the point at which one or more eigenvalues become complex is identified as the PT-symmetry breaking point or EP. Alternatively, the EP can be determined from the propagation pattern. Due to the exceptional growth that occurs after the EP, when eigenvalues turn complex, the propagation pattern exhibits a distinct difference before and after the EP. This approach provides a practical means of defining the EP.
A particularly intriguing application of PT-symmetry breaking points in photonics is in sensors, as systems near the EP exhibit a significantly enhanced response to even minor external disturbances. As the system approaches the n-th order Exceptional Point ($EP_{n}$), n eigenvalues and their corresponding eigenstates merge. The primary distinction between EPs and conventional degeneracies, such as Diabolic Points (DPs), lies in the enhanced sensitivity of the system to external disturbances [8].
Consequently, identifying the EP in optical waveguide arrays, and investigating the propagation patterns across the PT-symmetry breaking threshold is of critical importance. EPs play a fundamental role in PT-symmetric systems. As spectral singularities in open quantum systems, they exhibit an enhanced sensitivity to small external perturbations, making them promising candidates for advanced sensor technologies [8].
Thompson et al. [11] examined intensity and site-to-site correlations in disordered optical waveguide arrays under Anderson localization, using non-classical (squeezed) and classical-like (coherent) light. However, their system did not exhibit PT symmetry. In 2018, the PT-symmetry breaking point of two curved, closely spaced PT-symmetric waveguides, characterized by a small period, was numerically computed using COMSOL Multiphysics software [12]. The investigation of boundary effects on propagation dynamics and site-to-site correlations within optical waveguide arrays in the quantum regime was reported in 2023 [1]. In [10], the behavior of two-photon states in non-Hermitian systems across the EP was studied. The EP was identified by analyzing variations in the Hong-Ou-Mandel interference with two indistinguishable input photons. The objective of the present work is to compute the EP in a PT-symmetric optical waveguide array with a quantum input state. To this end, two methods are employed: one based on the analysis of the system’s eigenvalues, and the other through the examination of the propagation patterns.
The organization of the paper is as follows. In Section 1, the theoretical background is discussed in detail. Section 11 contains the numerical results and their analysis. A summary and concluding are given in Section 4.
2 Theory
The system being studied is an optical waveguide array containing N single-mode waveguides. This work focuses on analyzing the evolution of a quantum input state in a non-Hermitian PT-symmetric system. Therefore, to satisfy the PT-symmetry criteria, for an even number of waveguides (Figure 1(a)), the waveguides alternate between loss and gain. In addition, for an odd number of waveguides (Figure 1(b)), the central waveguide exhibits neither loss nor gain, while the waveguides on either side alternate between loss and gain.
Figure 1: PT-symmetric system for (a) even number, and (b) odd, number of waveguides
To analyze the propagation of a quantum input state through an optical waveguide array, it is essential to perform the calculations within the framework of quantum optics. Within this regime, the field amplitude is replaced by the creation and annihilation operators, which are used to describe the quantum properties of the field. Let $\hat{a}_{j}$ and $\hat{a}^{\dagger}_{j}$ denote the annihilation and creation operators for the j-th waveguide, respectively. The commutation relations can be expressed as follows [13]:
$[\hat{a}_{j},\hat{a}_{k}^{\dagger}]=\delta_{jk},\quad[\hat{a}_{j},\hat{a}_{k}]=0,\quad[\hat{a}_{j}^{\dagger},\hat{a}_{k}^{\dagger}]=0.$ (1)
If the propagation constant of the $j$-th waveguide is denoted by $\beta_{j}$ and the coupling coefficient between two adjacent waveguides is $c$ then the Hamiltonian describing the waveguide array is given by [13]:
$\hat{H}=\sum_{j}c\left(\hat{a}_{j+1}^{\dagger}\hat{a}_{j}+\hat{a}_{j}^{\dagger}\hat{a}_{j+1}\right)+\sum_{j}\beta_{j}\hat{a}_{j}^{\dagger}\hat{a}_{j},$ (2)
where the coupling between higher-order neighboring waveguides is neglected. The first term in equation (2) represents the transition of a photon from waveguide $j$ to $j+1$, while the second term corresponds to the reverse process. The final term represents the energy corresponding to each waveguide [13].
In equation (2), the propagation constant is given by $\beta_{j}=\beta_{0}-ig$ for waveguides exhibiting gain and by $\beta_{j}=\beta_{0}+ig$ for those with loss, where $g$ quantifies the gain and loss, and $\beta_{0}$ is the reference propagation constant that can be removed through normalization ($\beta_{0}=0$).
To analyze the evolution of the field operator $\hat{a}_{j}$, for the $j$-th waveguide, the Heisenberg equation is used [14]:
$i\frac{\partial\hat{a}_{j}}{\partial z}=-[\hat{H},\hat{a}_{j}].$ (3)
By substituting Hamiltonian (2) into equation (3), the resulting operator equation is obtained as follows:
$i\frac{\partial\hat{a}_{j}(z)}{\partial z}=c_{j-1}\hat{a}_{j-1}(z)+c_{j}\hat{a}_{j+1}(z)+\beta_{j}\hat{a}_{j}(z),$ (4)
where $j=1,2,\dots,N$. Since equation (4) is linear, there exists a linear relationship between the annihilation operator at propagation length $z$, denoted $\hat{a}_{j}(z)$, and the annihilation operator at the initial point ($z=0$), denoted $\hat{a}_{m}$ [11]:
$\hat{a}_{j}(z)=\sum_{m=1}^{N}G_{jm}(z)\hat{a}_{m}$ (5)
In this relation $G_{jm}(z)$, known as the Green’s function, connects the field at propagation length $z$ in the $j$-th waveguide to the initial field in the $m$-th waveguide. By substituting equation (5) into equation (4), the following equation is obtained to calculate $G_{jm}(z)$ ;
$i\frac{\partial G_{jm}(z)}{\partial z}$ $=c\,G_{j-1,m}(z)+c\,G_{j+1,m}(z)$
$\quad+\beta_{j}G_{j,m}(z)$ (6)
This system of equations consists of $N^{2}$ scalar differential equations that can be solved numerically to obtain the matrix elements of the Green’s function. According to equation (5), at $z=0$, the initial condition for solving these equations is $G_{jm}(z=0)=\delta_{jm}$.
To analyze the propagation pattern, one must calculate the intensity in each $j$-th waveguide as a function of the propagation distance $z$, which corresponds to the expectation value of the photon number in that waveguide.
$I_{j}(z)=\langle\hat{a}_{j}^{\dagger}(z)\hat{a}_{j}(z)\rangle$ (7)
Substituting equation (5) into equation (7) yields the following expression for the intensity:
$I_{j}(z)=\sum_{m,n}G_{jm}^{*}(z)\,G_{jn}(z)\,\langle\hat{a}_{m}^{\dagger}\hat{a}_{n}\rangle$ (8)
The resulting expression represents the probability of detecting photons in the $j$-th waveguide at a propagation distance of $z$.
The site-to-site correlation function is utilized to determine the probability of simultaneous photon detection in waveguides $i$ and $j$:
$g_{ji}^{(2)}(z)=\langle\hat{a}_{j}^{\dagger}(z)\,\hat{a}_{i}^{\dagger}(z)\,\hat{a}_{i}(z)\,\hat{a}_{j}(z)\rangle$ (9)
Substituting equation (5) into relation (9) yields the following expression for the site-to-site correlation:
$g_{ji}^{(2)}(z)$ $=\sum_{m,n,k,l}G_{jm}^{*}(z)\,G_{in}^{*}(z)$
×$\quad\times G_{ik}(z)\,G_{jl}(z)\,\langle\hat{a}_{m}^{\dagger}\hat{a}_{n}^{\dagger}\hat{a}_{k}\hat{a}_{l}\rangle$ (10)
Relations (8) and (10) show that the intensity and site-to-site correlation, can be calculated at any point along the propagation path if the Green’s function and initial state are known.
A primary aim of this work is to investigate how the propagation pattern and site-to-site correlations evolve for various quantum input states within the system. For this purpose, two categories of different input states have been investigated. The first category includes the number state $|m\rangle_{p}$, coherent state $|\alpha\rangle_{p}$ and squeezed state $|\xi\rangle_{p}$ input into a waveguide number $p$ (Please note that state $|\psi\rangle_{p}$ is a shorthand for exact state $|0\rangle_{1}|0\rangle_{2}\dots|\psi\rangle_{p}\dots|0\rangle_{N}$, where waveguides with zero input photons are not displayed). The second category is entangled NOON state input into two waveguides, $p$ and $q$, $\frac{|n\rangle_{p}|0\rangle_{q}+|0\rangle_{p}|n\rangle_{q}}{\sqrt{2}}$. In the latter scenario, two cases of the light input in the middle and boundary waveguides have been investigated. To this end, for different input states, the following expressions are utilized:
Number state $|n\rangle_{p}$,
$\langle\hat{a}_{m}^{\dagger}\hat{a}_{n}\rangle=n\,\delta_{np}\delta_{mp},$ (11)
$\langle\hat{a}_{m}^{\dagger}\hat{a}_{n}^{\dagger}\hat{a}_{k}\hat{a}_{l}\rangle=n(n-1)\,\delta_{mp}\delta_{np}\delta_{kp}\delta_{lp},$
$I_{j}(z)=n\,|G_{jp}(z)|^{2},$
$g_{ji}^{(2)}(z)=n(n-1)\,|G_{jp}(z)G_{ip}(z)|^{2}.$
Coherent state $|\alpha\rangle_{p}$,
$\langle\hat{a}_{m}^{\dagger}\hat{a}_{n}\rangle=|\alpha|^{2}\delta_{np}\delta_{mp},$ (12)
$\langle\hat{a}_{m}^{\dagger}\hat{a}_{n}^{\dagger}\hat{a}_{k}\hat{a}_{l}\rangle=|\alpha|^{4}\delta_{mp}\delta_{np}\delta_{kp}\delta_{lp},$
$I_{j}(z)=|\alpha|^{2}\,|G_{jp}(z)|^{2},$
$g_{ji}^{(2)}(z)=|\alpha|^{4}\,|G_{jp}(z)G_{ip}(z)|^{2}.$
Squeeze state $|\xi\rangle_{p}$,
$\langle\hat{a}_{m}^{\dagger}\hat{a}_{n}\rangle=\sinh^{2}(\xi)\,\delta_{np}\delta_{mp},$ (13)
$\langle\hat{a}_{m}^{\dagger}\hat{a}_{n}^{\dagger}\hat{a}_{k}\hat{a}_{l}\rangle=\sinh^{2}(\xi)\left(\sinh^{2}(\xi)+2\right)\delta_{mp}\delta_{np}\delta_{kp}\delta_{lp},$
$I_{j}(z)=\sinh^{2}(\xi)\,|G_{jp}(z)|^{2},$
$g_{ji}^{(2)}(z)=\sinh^{2}(\xi)\left(\sinh^{2}(\xi)+2\right)|G_{jp}(z)G_{ip}(z)|^{2}.$
Entangled NOON state,
$\begin{split}&\langle\hat{a}_{m}^{\dagger}\hat{a}_{n}\rangle=\frac{n}{2}\big{(}\delta_{np}\delta_{mp}+\delta_{nq}\delta_{mp}+\delta_{np}\delta_{mq}+\delta_{nq}\delta_{mq}\big{)},\\[6.0pt] &\langle\hat{a}_{m}^{\dagger}\hat{a}_{n}^{\dagger}\hat{a}_{k}\hat{a}_{l}\rangle=\frac{n(n{-}1)}{2}\bigg{(}\delta_{mp}\delta_{np}\delta_{kp}\delta_{lp}+\delta_{mq}\delta_{nq}\delta_{kq}\delta_{lq}\\ &\quad+\delta_{n,2}\big{(}\delta_{mp}\delta_{np}\delta_{kq}\delta_{lq}+\delta_{mq}\delta_{nq}\delta_{kp}\delta_{lp}\big{)}\bigg{)},\\[6.0pt] &I_{j}(z)=\frac{n}{2}\Big{(}|G_{jp}(z)|^{2}+|G_{jq}(z)|^{2}\\ &\quad+G^{*}_{jp}(z)\,G_{jq}(z)+G_{jp}(z)\,G^{*}_{jq}(z)\Big{)},\\[6.0pt] &g_{ji}^{(2)}(z)=\frac{n(n{-}1)}{2}\Big{(}|G_{jp}(z)G_{ip}(z)|^{2}+|G_{jq}(z)G_{iq}(z)|^{2}\\ &\quad+\delta_{n,2}\big{(}G^{*}_{jp}(z)G^{*}_{ip}(z)G_{jq}(z)G_{iq}(z)\\ &\qquad+G^{*}_{jq}(z)G^{*}_{iq}(z)G_{jp}(z)G_{ip}(z)\big{)}\Big{)}\end{split}$ (14)
An additional objective of this study is to determine the EP ($g_{c}$) of the system. To this end, two different approaches are evaluated. Since the propagation pattern exhibits oscillatory behavior for $g<g_{c}$ and transitions to an exponential energy growth (in the linear regime) for $g>g_{c}$, the propagation pattern itself can be used to identify the exceptional point. For this purpose, the following criterion is defined:
$I_{r}=\exp\left(-\alpha\left|\frac{\bar{I}_{2}-\bar{I}_{1}}{\bar{I}_{2}+\bar{I}_{1}}\right|^{\beta}\right)$ (15)
The determination is based on the spatial intensity distribution along the propagation length, with $\bar{I}_{1}$ and $\bar{I}_{2}$ representing the mean intensities over all waveguides in the first half ($0<z<\frac{z_{\text{max}}}{2}$) and second half ($\frac{z_{\text{max}}}{2}<z<z_{\text{max}}$) of the propagation length $z_{\text{max}}$, respectively. Besides, the parameters $\alpha$ and $\beta$ are two positive constants chosen to exhibit a significant change at the EP. In this study, the values $\alpha=\beta=10$ are employed. Before the EP, the energy exhibits oscillatory behavior, so $\bar{I}_{1}\simeq\bar{I}_{2}$. In this case, $\left|\frac{\bar{I}_{2}-\bar{I}_{1}}{\bar{I}_{2}+\bar{I}_{1}}\right|^{\beta}$ is approximately zero, which results in $I_{r}=1$. However, after the EP, as the propagation length increases, the energy grows exponentially, so $\bar{I}_{1}\ll\bar{I}_{2}$. Therefore, $\left|\frac{\bar{I}_{2}-\bar{I}_{1}}{\bar{I}_{2}+\bar{I}_{1}}\right|^{\beta}$ approaches one, and $I_{r}=0$. Subsequently, by plotting $I_{r}$ versus gain/loss factor $g$ and investigating the transition from one to zero in the $I_{r}$ criterion, the PT-symmetry breaking point $g_{c}$ can be identified.
A more accurate approach to locate the EP is to diagonalize the system Hamiltonian and calculate its eigenvalues. The EP is found at a critical gain/loss strength $g_{c}$ where the Hamiltonian’s eigenvalues shift from real to complex, which can be observed by plotting their imaginary parts versus $g$. The Hamiltonian (2) is defined in an infinite-dimensional Hilbert space. Consequently, determining its eigenvalues presents a significant analytical and computational challenge. Because the Hamiltonian (2) preserves the total photon number, as ensured by the commutation relation $[\hat{H},\hat{n}_{t}]=0$ with $\hat{n}_{t}=\sum_{j}\hat{a}_{j}^{\dagger}\hat{a}_{j}$, the system’s dynamics can be analyzed within separate photon-number subspaces (zero, one, two photons, etc.), simplifying the problem.
In general, for a waveguide array consisting of $N$ waveguides, the subspace corresponding to $m$ photons forms a Hilbert space of dimension $d=\binom{m+N-1}{m}.$ The standard basis of this subspace is given by the state vector $|n_{1},n_{2},n_{3},\dots,n_{N}\rangle=|n_{1}\rangle_{1}|n_{2}\rangle_{2}\dots|n_{N}\rangle_{N}$, subject to the constraints $n_{j}\geq 0$ and $\sum_{j=1}^{N}n_{j}=m$. This state stands for a case with $n_{j}$ photons at j-th waveguide ($j=1,2,..,N$). In order to calculate the eigenvalues within this subspace, it is necessary to first determine the elements of the Hamiltonian matrix in the specified basis. Thus, we obtain:
$\hat{H}\left|\dots,n_{j},n_{j+1},\dots\right\rangle=\sum_{j=1}^{N}\beta_{j}n_{j}\left|\dots,n_{j},n_{j+1},\dots\right\rangle\\ +c\sum_{j=1}^{N-1}\Big{\{}\sqrt{(n_{j}+1)n_{j+1}}\left|\dots,n_{j}{+}1,n_{j+1}{-}1,\dots\right\rangle\\ +\sqrt{n_{j}(n_{j+1}+1)}\left|\dots,n_{j}{-}1,n_{j+1}{+}1,\dots\right\rangle\Big{\}}$ (16)
Therefore, the matrix elements of the Hamiltonian will be as follows:
$\langle n_{1}^{\prime},n_{2}^{\prime},n_{3}^{\prime},\dots,n_{N}^{\prime}|\hat{H}|n_{1},n_{2},n_{3},\dots,n_{N}\rangle$ $=$ (17)
$\left(\sum_{j=1}^{N}\beta_{j}n_{j}^{\prime}\right)\delta_{n_{1}^{\prime},n_{1}}\delta_{n_{2}^{\prime},n_{2}}\dots\delta_{n_{N}^{\prime},n_{N}}$
$+c\sum_{j=1}^{N}\left\{\sqrt{n_{j}^{\prime}(n_{j+1}^{\prime}+1)}\delta_{n_{1}^{\prime},n_{1}}\delta_{n_{2}^{\prime},n_{2}}\dots\delta_{n_{j}^{\prime},n_{j}+1}\right.$
$\left.\delta_{n_{j+1}^{\prime},n_{j+1}-1}\dots\delta_{n_{N}^{\prime},n_{N}}\right.$
$+\sqrt{(n_{j}^{\prime}+1)n_{j+1}^{\prime}}\delta_{n_{1}^{\prime},n_{1}}\delta_{n_{2}^{\prime},n_{2}}\dots\delta_{n_{j}^{\prime},n_{j}-1}$
$\left.\delta_{n_{j+1}^{\prime},n_{j+1}+1}\dots\delta_{n_{N}^{\prime},n_{N}}\right\}$
Equation (17) can be used to calculate the eigenvalues of the Hamiltonian in the $m$-photon subspace and to examine their imaginary parts to find the EP. In the next sections, we will calculate and compare the EP of a waveguide array using the two methods discussed.
3 Numerical Results
Numerical results for the propagation pattern, site-to-site correlation, and exceptional point are presented in this section. As a result, the propagation length is normalized with respect to the coupling constant $c$ ($c=1$). First, the propagation pattern is investigated for a waveguide array consisting of $N=50$ waveguides, with quantum input states including separable and entangled states, in both Hermitian and non-Hermitian systems. Figure 2 shows the results in the Hermitian regime ($g=0$), for different quantum input states injected into waveguide $p=25$, including the number state $|n\rangle_{p}$, coherent state $|\alpha\rangle_{p}$, and squeezed state $|\xi\rangle_{p}$, each prepared with an average of one photon ($n=1$, $\alpha^{2}=1$, $\sinh^{2}\xi=1$). Additionally, two types of entangled states are considered: $\frac{\left(|0\rangle_{1}|1\rangle_{2}+|1\rangle_{1}|0\rangle_{2}\right)}{\sqrt{2}}$ and $\frac{\left(|0\rangle_{24}|1\rangle_{25}+|1\rangle_{24}|0\rangle_{25}\right)}{\sqrt{2}}$, which are injected into the 1st–2nd and 24th–25th waveguides, respectively. Since the average number of input photons in the separable states is identical, the resulting propagation patterns are equivalent for all three cases, consistent with equations (11) to (13).
Figure 2: Propagation pattern of a waveguide array with $N=50$ waveguides in the Hermitian regime ($g=0$) for: (a) number state $|n\rangle_{p}$ with $n=1$, (b) coherent state $|\alpha\rangle_{p}$ with $\alpha^{2}=1$, (c) squeezed state $|\xi\rangle_{p}$ with $\sinh^{2}(\xi)=1$, (d) entangled state between the 1st and 2nd waveguides, and (e) entangled state between the 24th and 25th waveguides.
Moreover, the figures distinctly exhibit the characteristic features of discrete diffraction and highlight the impact of boundary reflections. This figure demonstrates the propagation pattern within the Hermitian regime, characterized by the absence of gain or loss ($g=0$). Upon the introduction of gain or loss to the system, and when its magnitude remains below the critical value $g_{c}$, the system stays in the PT-symmetric regime. Figure (3) illustrates that in this regime ($g=0.03$), the propagation pattern is nearly identical to that in the Hermitian regime. However, unlike in the Hermitian state, the total intensity of light in the PT-symmetric phase is not fixed, instead exhibiting oscillations throughout the propagation due to the effects of gain and loss. For better comparison, in all plots related to the propagation pattern, the total intensity is normalized at each propagation distance.
Figure 3: Propagation pattern of a waveguide array with $N=50$ waveguides in the PT-symmetry regime ($g=0.03$) for: (a) number state $|n\rangle_{p}$ with $n=1$, (b) coherent state $|\alpha\rangle_{p}$ with $\alpha^{2}=1$, (c) squeezed state $|\xi\rangle_{p}$ with $\sinh^{2}(\xi)=1$, (d) entangled state between the 1st and 2nd waveguides, and (e) entangled state between the 24th and 25th waveguides.
In Figure (4), the propagation pattern of a waveguide array consisting of $N=50$ waveguides with a quantum input state is investigated in the broken PT-symmetry regime $g>g_{c}$ ($g=0.6$). As energy grows exponentially in this regime, the waveguides with gain near the input one exhibit much higher intensity than others. Therefore, after normalization, the intensity contributions from the rest of the waveguides are negligible and nearly vanish in the propagation pattern.
Figure 4: Propagation pattern of a waveguide array with $N=50$ waveguides in the broken PT-symmetry regime ($g=0.6$) for: (a) number state $|n\rangle_{p}$ with $n=1$, (b) coherent state $|\alpha\rangle_{p}$ with $\alpha^{2}=1$, (c) squeezed state $|\xi\rangle_{p}$ with $\sinh^{2}(\xi)=1$, (d) entangled state between the 1st and 2nd waveguides, and (e) entangled state between the 24th and 25th waveguides.
The study of site-to-site correlations for input number state $|n\rangle_{p}$ (Eq. (11)) and NOON state (Eq. (14)) requires considering input states with more than one photon, as $g_{ji}^{(2)}$ quantifies the probability of simultaneous detection of two photons in distinct waveguides $i$ and $j$. Accordingly, in this work, we compute the site-to-site correlation for different input states with average photon number of two (relations  (11) to (14)). Figure (5) shows the site-to-site correlation for a waveguide array with $N=50$ waveguides at a propagation distance of $z=10$ in the Hermitian regime, where $g=0$.
Apart from a factor, the site-to-site correlation is identical for the number state, coherent state, and squeezed state, with the highest values concentrated near the four corners. This result demonstrates that, in the given state and propagation pattern at $z=10$, the probability of detecting two photons in close proximity (either near waveguide 7 or 43) is approximately identical to the probability of detecting two photons at a specific distance — one near waveguide 7 and the other near waveguide 43.
Figure 5: Site-to-site correlation a waveguide array with $N=50$ waveguides in the Hermitian regime ($g=0$) for: (a) number state $|n\rangle_{p}$ with $n=2$, (b) coherent state $|\alpha\rangle_{p}$ with $\alpha^{2}=2$, (c) squeezed state $|\xi\rangle_{p}$ with $\sinh^{2}(\xi)=2$, (d) entangled state between the 1st and 2nd waveguides, and (e) entangled state between the 24th and 25th waveguides at propagation distance $z=10$.
The site-to-site correlation shown in Figure 5(e) for an entangled state introduced at the middle of the array indicates that the photons are most probably detected at widely separated positions, approximately at waveguides 7 and 43. This pattern is a signature of spatial anti-bunching. In comparison, Figure 5(d) highlights that entangled photons entering through the array’s boundaries tend to remain spatially close, exhibiting a peak near waveguide 20. Therefore, this plot shows spatial photon bunching because the photons exit around a specific waveguide.
Similar to Figure (5), Figure (6) illustrates the case of the unbroken PT-symmetric phase, where the gain/loss parameter satisfies $g<g_{c}$ ($g=0.03$). By comparing Figure (5) and Figure (6), it is evident that the statistical features of the PT-symmetric phase, such as intensity distribution and site-to-site correlation, closely resemble those observed in the Hermitian regime. In other words, the site-to-site correlation exhibits similar behavior regardless of whether the system operates in the PT-symmetric or Hermitian regime. Due to the exact balance between gain and loss, the optical system in the PT-symmetric regime maintains a stable evolution, analogous to the Hermitian case where neither gain nor loss is present.
Figure 6: Site-to-site correlation a waveguide array with $N=50$ waveguides in the PT-symmetry regime ($g=0.03$) for: (a) number state $|n\rangle_{p}$ with $n=2$, (b) coherent state $|\alpha\rangle_{p}$ with $\alpha^{2}=2$, (c) squeezed state $|\xi\rangle_{p}$ with $\sinh^{2}(\xi)=2$, (d) entangled state between the 1st and 2nd waveguides, and (e) entangled state between the 24th and 25th waveguides at propagation distance $z=10$.
In Figure (7), the site-to-site correlation functions are shown for different quantum input states under the broken PT-symmetry condition, characterized by $g>g_{c}$ ($g=0.6$). It is demonstrated that, in the broken PT-symmetric regime, the correlation behavior is modified in contrast to the Hermitian and unbroken PT-symmetric cases. In this regime, the value of the correlation function reaches its maximum near the input waveguide, indicating that the two photons are most likely to be detected in the nearby gain waveguides.
As has been observed thus far, the system exhibits distinct behavior in the PT-symmetric and broken PT-symmetric regimes. Therefore, calculating the EP is essential to more accurately understand the system’s behavior.
Figure 7: Site-to-site correlation a waveguide array with $N=50$ waveguides in the Broken PT-symmetry regime (g=0.6) for: (a) number state $|n\rangle_{p}$ with $n=2$, (b) coherent state $|\alpha\rangle_{p}$ with $\alpha^{2}=2$, (c) squeezed state $|\xi\rangle_{p}$ with $\sinh^{2}(\xi)=2$, (d) entangled state between the 1st and 2nd waveguides, and (e) entangled state between the 24th and 25th waveguides at propagation distance $z=10$.
Thus, based on the calculated propagation patterns for different quantum input states, the EP has been determined using the previously introduced criterion $I_{r}$. Figure (8) illustrates the variation of $I_{r}$ with respect to the gain/loss coefficient $g$, for an entangled input state containing one photon distributed across the 24th and 25th waveguides in an array of $N=10$ waveguides.
The EP in this case is the transition point from one to zero in $I_{r}$ which is $g_{c}\simeq 0.289$ for this waveguide array. Through numerical analysis, it has been confirmed that the EP remains invariant with respect to the choice of input state; That is, the position of the EP remains unchanged for different quantum input states. This is essential because $g_{c}$ is a characteristic of the system itself and its corresponding Hamiltonian, not its input or initial state.
Figure 8: $I_{r}$ as a function of gain/loss parameter $g$ for an entangled input state in the 24th–25th waveguides of a $N=10$ waveguide array.
The EP is typically identified as the critical value $g_{c}$ at which the system’s eigenvalues become complex. As described in the theoretical section, this is done by constructing the reduced Hamiltonian in the $m$-photon subspace and computing its eigenvalues. The EP is then located by plotting the maximum imaginary part of the eigenvalues versus the gain/loss parameter $g$. Figure (9) shows the maximum imaginary part of all eigenvalues as a function of $g$ for a waveguide array with $N=10$ in the single-photon subspace ($m=1$).
Figure 9: Maximum imaginary part of the eigenvalues as a function of the gain/loss parameter $g$ for a waveguide array with $N=10$ in the single-photon ($m=1$) subspace.
Figure 10: Imaginary parts of the eigenvalues as a function of the gain/loss parameter $g$ for a waveguide array with $N=10$, shown for different photon-number subspaces.
Based on the plot, the EP for this case is determined to be $g_{c}=0.284$, with a difference of only 0.005 compared to the EP obtained using $I_{r}$ (Figure (8)). This result shows the close values obtained by both methods.
The maximum imaginary part of the eigenvalues, as observed for a waveguide array with $N=10$ waveguides, in subspaces including one to five photons, is presented in figure (10). As shown in figure (10), although the maximum value of the imaginary parts of the eigenvalues differs across various photon subspaces, the onset of complex eigenvalues does not depend on the photon number $m$ in the subspace. In other words, increasing the number of photons does not affect the value of the EP. Therefore, it can be concluded that the method of calculating the EP using the reduced Hamiltonian in the $m$-photon subspace provides a general approach, and the EP of the full Hamiltonian is well-obtained.
Afterwards, in addition to comparing the exceptional point $g_{c}$ determined by the two methods, the influence of the number of waveguides on the EP is investigated. Since the PT-symmetry pattern differs between arrays with even and odd numbers of waveguides—where even-numbered arrays exhibit alternating gain and loss, while in odd-numbered arrays the central waveguide is neutral and the rest alternate in gain and loss, the dependence of the EP on the number of waveguides is investigated separately for the even and odd cases.
Figure 11: Comparison of the EP values obtained using both methods for (a) even and (b) odd numbers of waveguides.
As shown in Figure (11), an increase in the number of waveguides leads to a decrease in $g_{c}$, indicating that the system enters the broken $\mathcal{PT}$-symmetric regime at lower gain/loss values. Furthermore, Figure (11) confirms the consistency of the two approaches for identifying the $\mathcal{PT}$-symmetry breaking threshold. Another notable observation is that the value of the EP in arrays with an odd number of waveguides is higher than in those with an even number. This may be attributed to the presence of a central waveguide without gain or loss in arrays with an odd number of waveguides. The consistency between the EPs obtained via the two approaches is considered highly important. Given that the diagonalization of the Hamiltonian in PT-symmetric systems—especially in high-dimensional Hilbert spaces—is typically time-consuming, the $I_{r}$-based approach, which utilizes the propagation pattern, can serve as an effective alternative to the direct computation of eigenvalues.
4 Summary and Conclusion
The propagation patterns in Hermitian and non-Hermitian (PT-symmetric) systems have been analyzed for number states, coherent states, squeezed states, and entangled states, and it was found that entangled states exhibit a propagation behavior distinct from the other types of quantum states. Furthermore, the site-to-site correlation function was investigated for different quantum input states. The second-order correlation function indicates the probability of simultaneously detecting two photons in distinct waveguides. The results indicate that, in the unbroken PT-symmetric regime, the statistical properties associated with photon intensity and site-to-site correlations remain stationary due to the balanced gain and loss, resembling those of a Hermitian system without gain or loss. One of the most notable differences in the site-to-site correlation pattern in the broken PT-symmetric phase is the concentration of the correlation function around the input waveguide. In other words, in this regime, photons are most likely to be detected in the gain-enhanced waveguides near the injection site.
In addition, the EP was determined using two distinct methods. The results demonstrated consistency between the two approaches. Although computing the eigenvalues of the system is the direct approach for determining the EP, the strong agreement between both methods suggests that the $I_{r}$ criterion, which is based on the propagation pattern, can serve as an effective alternative for identifying the PT-symmetry breaking point. Additionally, the variation of the EP with respect to the number of waveguides was investigated, revealing that the EP value decreases as the number of waveguides in the array increases. In other words, the system enters the broken $\mathcal{PT}$-symmetric phase at a lower gain/loss threshold.